OpenStudy (anonymous):

5x-3x-9=5x does x=12 or x=-3

OpenStudy (anonymous):

Add -5x to each side. -3x-9 = 0 x = -3

OpenStudy (madhu.mukherjee.946):

x = -3

OpenStudy (anonymous):

5x−3x−9=5x Simplify 5x−3x−9=5x 5x+−3x+−9=5x (5x+−3x)+(−9)=5x(Combine Like Terms) 2x+−9=5x 2x−9=5x Subtract 5x from both sides. 2x−9−5x=5x−5x −3x−9=0 Add 9 to both sides. −3x−9+9=0+9 −3x=9 Divide both sides by -3. −3x/−3 = 9/−3 x=−3